# The Anderson transition – a view from Krylov space

### v1

@@ III.1 Weak Disorder
Figure 4: Zero mode probability density in Krylov space for $W=4$. The dot-dashed line represents the geometric average of the zero-modes constructed in each sample whereas the colored lines are obtained from the sample-averaged $b_{n}$: the two are in complete agreement. There is an extremely small exponential decay which vanishes as a function of increasing $L$. The inset shows the disorder averaged dimerization of the Lanczos coefficients for different system sizes $L=14,18,22,23$, revealing its inverse scaling with the volume, roughly as $\simeq 6.3L^{-3}$. There are $N_{s}=1000$ samples for all sizes except $L=23$ for which there are $N_{s}=500$.
In the extended regime, we’d expect the zero-mode to be delocalized in Krylov space. However, careful analysis of this zero-mode, shown in Fig. 4, reveals that the zero-mode is actually exponentially localized albeit extremely weakly. While this might at first sight seem to contradict our earlier statements, it’s perfectly consistent. Recall that exponential localization is associated with perfect dimerization of the Lanczos coefficients and hence with a spectral gap. Since the numerics is performed on a finite size system $L$ and we’re going to Krylov space indices which far exceed the linear system size, we’re actually picking up on the finite size gap. On the delocalized side, the system is chaotic and thus exhibits level repulsion. As such, the typical smallest level spacing is of $O(1/V)$. As shown in Fig. 4, the average dimerization, and thus the associated Krylov localization length, goes to zero in the thermodynamic limit as $1/L^{3}$. This is confirmed with our fit of the coefficients, in the inset of Fig. 3 where $\epsilon\approx 1/L^{3}\approx 9\times 10^{-5}$.
While this makes sense from the perspective of the spectral finite-size effects, note that the total size of Krylov space is quadratically larger $D_{K}=O(V^{2})$. Over the entire range from $O(V)$ to $O(V^{2})$ the zero-mode is thus exponentially small. That this is inevitable also follows from a simple mathematical constraint. Consider again the two-time correlation function given by expression (3). In Krylov space, the long time limit of the correlation function is simply given by the zero-mode probability on the first site:
@@
where $\left|n\right$> denote the eigenstates of the many-body Hamiltonian, i.e. $H\left|n\right>=E_{n}\left|n\right$>. Both expressions should result in the same answer, consequently equation (13) constraints $|Q_{1}|^{2}$. For a system with conserved particle number, that is otherwise ergodic, one would expect (13) to be of $O(1/V)$. The latter explains why the zero-mode needs to be localized in Krylov space on the first $O(V)$ sites. At present, we can be slightly more specific. Since our theory is non-interacting, we can recast (13) in terms of single particle eigenstates $\psi_{n}(x)$. Recall our initial operator $O_{1}=c^{\dagger}_{x}c_{x}-1/2$, such that one arrives at:
@@ III.2 Critical Disorder
Figure 5: The zero-mode probability density close to the critical point ($W=16$) is shown to decay as a powerlaw $n^{\alpha}$ with an exponent $\alpha$ approaching $\alpha\approx-0.58$ (shown by the grey dashed line). The dot-dashed line represents the geometric average of the zero-modes constructed in each sample whereas the colored lines are obtained from the sample-averaged $b_{n}$: the two are in complete agreement.
The Krylov space properties of the critical point can be constrained by the same analysis we’ve conducted on the delocalized side. The Anderson critical point is known to exhibit multi-fractality [94, 4] and was recently studied in detail in Ref. [95]. This multi-fractality manifests itself as powerlaw scaling of the survival probability $C(t)\sim t^{-\beta}$, with $\beta\approx 0.42$. At the critical point, the Thouless time approaches the Heisenberg time (which is $O(1/V)$, implying
@@
Consequently, the zero mode must itself decay like a powerlaw:
@@
Since the finite size gap, i.e. the inverse Heisenberg time, is still $O(1/V)$, one immediately finds that $\beta=1-2\alpha$ and $\alpha\approx-0.29$. This scaling is indeed numerically observed, as shown in Fig. 5, where for increasing $L$ the zero-mode is approaching the desired power-law.
These conclusions follow from rather general physical constraints. At this point, it’s interesting to tie this back to the behavior of the Lanczos coefficients. A powerlaw decaying zero mode implies (c.f. eq. (7)) that $\kappa(n)\sim\alpha\log(n)$. Consequently, the Lanczos coefficients asymptotically behave like:
@@
This is consistent with the transition point of our model: $\gamma=0$. If the dimerization would decay any faster than $1/n$, it would result in delocalization, whereas faster decay would result in a localized zero-mode.
@@ III.3 Strong Disorder
Figure 6: Zero-mode probability density for the three-dimensional Anderson model in Krylov space over a wide range of the disorder $W$. The results are shown as a function of $n^{1/2d}$ (with $d=3$), the anticipated stretched coordinate deep in the localized phase. The dot-dashed line represents the geometric average of the zero-modes constructed in each sample whereas the colored lines are obtained from the sample-averaged $b_{n}$: the two are in complete agreement.
@@
Figure 7: The zero-mode probability density in Krylov space for Anderson models in different spatial dimensions. The disorder is chosen to be much greater than the critical disorder such that all states are well localized (note the critical disorder in four dimensions is estimated to be $W_{c}\approx 33.2$ [96]). The results are shown as a function of the anticipated stretched Krylov coordinate $n^{1/2d}$, rescaled by the maximum value $n_{\text{max}}^{\frac{1}{2d}}$ so that all curves appear in the same scale. $d=1$ appears in the inset which has the same axes but an extended vertical range. The dot-dashed line represents the geometric average of the zero-modes constructed in each sample whereas the colored lines are obtained from the sample-averaged $b_{n}$: the two are in complete agreement. Results for spatial dimensions $d=1,2,3,4$ have disorders $W=10,20,40,80$, lattice lengths $L=12000,110,23,10$, and number of samples $N_{s}=500,1000,1000,1000$ respectively.
As such, we suggest that $\gamma$ is fixed to some value $0<\gamma<1$ deep in the localized phase. Recall that, both at the critical point and in the delocalized phase, we mapped real space properties to Krylov space by some rather generic arguments. Proceeding in the same vein, the absence of level repulsion in the delocalized phase suggests that the zero mode can now span the entire Krylov space of size $O(V^{2})$. At the same time, the only relevant length scale is the localization length $\xi$, implying the only dimensionless parameter is $L/\xi$. As such, one simply expects Krylov space to be some rescaled version of real space, i.e. $n]^{1/2d}=L$ where $d$ is the number of spatial dimensions of the model.
@@ IV Conclusions and Outlook
In summary, we have described the form of LIOMs in the Anderson model in Krylov space as generated by projectors on single sites. There are three types of behaviors – stretched exponential decay inside the localized phase (with geometric exponent $\gamma=1/2d$ set by real space dimension), delocalization on real-space volume $V=\sqrt{D_{K}}$ states in the metallic phase, and powerlaw decay at the critical point. Significantly, these results may be interpreted using effective disorder-free chains obtained from disorder averaged Krylov coefficients, which allows considerable further insight into the Anderson transition: e.g. the significance of Krylov dimerization as a mechanism of localizing operators despite having a gapless Liouvillian.
Our work may be extended to other aspects of the Anderson transition, for example to perhaps enable the tracking of the mobility edge as a function of typical energy. It also sets up a framework for extending these results to interacting models where the coefficients grow with some power $b_{n}\sim n^{\delta}$ (with $\delta\leq 1$). If the dimerization decays like a powerlaw as in our model $\sim n^{-(1-\gamma)}$ then – since only the ratio of Lanczos coefficients matters for the zero-mode – the growth simply re-normalizes the power to be $\sim n^{-(1-\gamma-\delta)}$. Consequently, the faster the coefficients grow, i.e. the faster the operator growth, the slower the dimerization must decay in order for conserved quantities to exist. For spin systems debated to show MBL, such as the disordered Heisenberg model, the coefficients have been found to grow as $\sim n/\log{n}$ on average, with some dimerization [80, 97]. The $\log$ correction would only be relevant at the critical power implying that the dimerization has to be constant for conserved quantities to exist (i.e. $\gamma=1$, its maximum allowed value). Additionally, it would be interesting to understand how the dimerization behaves in interacting integrable models such as the XXZ-chain, where there is typically square root growth ($\delta=0.5$) and thus only $1\geq\gamma\geq 0.5$ is allowed to host the conserved quantities we know to exist.

### v2

@@ III.1 Weak Disorder
Figure 4: Zero-mode probability density in Krylov space for $W=4$ in the extended regime. The dot-dashed line represents the geometric average of the zero-modes constructed in each sample whereas the colored lines are obtained from the sample-averaged $b_{n}$: the two are in complete agreement. There is an extremely small exponential decay which vanishes as a function of increasing $L$. The inset shows the disorder averaged dimerization of the Lanczos coefficients for different system sizes $L=14,18,22,23$, revealing its inverse scaling with the volume roughly as $\simeq 6.3L^{-3}$. There are $N_{s}=1000$ samples for all sizes except $L=23$ for which there are $N_{s}=500$.
In the extended regime we might expect the zero-mode to be delocalized in Krylov space. However, careful analysis of this zero-mode, shown in Fig. 4, reveals that the zero-mode is actually exponentially localized albeit extremely weakly. While this might at first sight seem to contradict our earlier statements, it’s perfectly consistent. Recall that exponential localization is associated with perfect dimerization of the Lanczos coefficients and hence with a spectral gap. Since the numerics are performed on a finite size system $L$ and we reach Krylov space indices which far exceed this linear system size, we are actually picking up on the finite size gap. On the delocalized side, the system is chaotic and thus exhibits level repulsion. As such, the typical smallest level spacing is of $O(1/V)$. As shown in Fig. 4 the average dimerization, and thus the associated Krylov localization length, goes to zero in the thermodynamic limit as $1/L^{3}$. This is confirmed with our fit of expression 10 in the inset of Fig. 3 where $\epsilon\approx 1/L^{3}\approx 9\times 10^{-5}$.
While this makes sense from the perspective of the spectral finite-size effects, note that the total size of Krylov space is quadratically larger $D_{K}=O(V^{2})$. Over the entire range from $O(V)$ to $O(V^{2})$ the zero-mode is thus exponentially small. That this is inevitable also follows from a simple mathematical constraint. Consider again the two-time correlation function given by expression (3). In Krylov space, the long time limit of the correlation function in expression (9) is simply given by the zero-mode probability on the first site:
@@
where $\left|n\right$> denote the eigenstates of the many-body Hamiltonian, i.e. $H\left|n\right>=E_{n}\left|n\right$>. Both expressions should result in the same answer, and consequently expression (12) constrains $|Q_{1}|^{2}$. For a system with conserved particle number that is otherwise ergodic, one would expect (12) to be of $O(1/V)$. The latter explains why the zero-mode needs to be localized in Krylov space on the first $O(V)$ sites. Now we can be slightly more specific. Since our theory is non-interacting, we can recast (12) in terms of single particle eigenstates $\psi_{n}(x)$. Recall our initial operator $O_{1}=c^{\dagger}_{x}c_{x}-1/2$, such that one arrives at:
@@ III.2 Critical Disorder
Figure 5: The zero-mode probability density for the three-dimensional Anderson model in Krylov space close to the critical point ($W=16$) is shown to decay as a powerlaw $n^{-2\alpha}$ with an exponent $\alpha$ approaching $\alpha\approx 0.29$ (shown by the dotted line). The dot-dashed line represents the geometric average of the zero-modes constructed in each sample whereas the colored lines are obtained from the sample-averaged $b_{n}$: the two are in complete agreement.
The Krylov space properties of the critical point can be constrained by the same analysis we’ve conducted on the delocalized side. The Anderson critical point is known to exhibit multi-fractality [Janssen1994_Multifractal_localization, Evers_2008_AndersonTransitions] and was recently studied in detail in Ref. [Hopjan2023_Scale_InvariantSurvivalProb]. This multi-fractality manifests itself as powerlaw scaling of the survival probability $C(t)\sim t^{-\beta}$, with $\beta\approx 0.42$. At the critical point, the Thouless time approaches the Heisenberg time (which is $O(1/V)$, implying
@@
Consequently, the zero-mode must itself decay like a powerlaw:
@@
Since the finite size gap, i.e. the inverse Heisenberg time, is still $O(1/V)$, one immediately finds that $\beta=1-2\alpha$ and $\alpha\approx 0.29$. This scaling is indeed numerically observed, as shown in Fig. 5, where for increasing $L$ the zero-mode is approaching the expected power-law.
These conclusions follow from rather general physical constraints. At this point, it’s interesting to tie this back to the behavior of the Lanczos coefficients. A powerlaw decaying zero-mode implies (c.f. expression (7)) that $\kappa(n)\sim\alpha\log(n)$. Consequently, the Lanczos coefficients asymptotically behave like:
@@
This is consistent with the transition point of our model: $\gamma=0$. If the dimerization would decay any faster than $1/n$, it would result in delocalization, whereas slower decay would result in a localized zero-mode.
@@ III.3 Strong Disorder
Figure 6: Zero-mode probability density for the three-dimensional Anderson model in Krylov space approaches a stretched exponential decay as $W$ is increased into the localized regime ($W>16$). The results are shown as a function of $n^{1/2d}$ (with $d=3$), the predicted stretching coordinate deep in the localized phase. The dot-dashed line represents the geometric average of the zero-modes constructed in each sample whereas the colored lines are obtained from the sample-averaged $b_{n}$: the two are in complete agreement.
@@
Figure 7: The zero-mode probability density in Krylov space for Anderson models in different spatial dimensions $d$. In all cases the disorder is chosen to be much greater than the critical disorder such that all states are well localized (note the critical disorder in four dimensions is estimated to be $W_{c}\approx 33.2$ [PMarkos_1994_Metal_InsulatorTransition4D]). The results are shown as a function of the anticipated stretched Krylov coordinate $n^{1/2d}$, rescaled by the maximum value $n_{\text{max}}^{\frac{1}{2d}}$ so that all curves appear in the same scale. $d=1$ appears in the inset which has the same axes but an extended vertical range. The dot-dashed line represents the geometric average of the zero-modes constructed in each sample whereas the colored lines are obtained from the sample-averaged $b_{n}$: the two are in complete agreement. Results for spatial dimensions $d=1,2,3,4$ have disorders $W=10,20,40,80$, lattice lengths $L=12000,110,23,10$, and number of samples $N_{s}=500,1000,1000,1000$ respectively.
As such, we suggest that $\gamma$ is fixed to some value $0<\gamma<1$ deep in the localized phase. Recall that, both at the critical point and in the delocalized phase, we mapped real space properties to Krylov space by some rather generic arguments. Proceeding in the same vein, the absence of level repulsion in the delocalized phase suggests that the zero-mode can now span the entire Krylov space of size $O(V^{2})$. At the same time, the only relevant length scale is the localization length $\xi$, implying the only dimensionless parameter is $L/\xi$. As such, one simply expects Krylov space to be some rescaled version of real space, i.e. $n]^{1/2d}=L$ where $d$ is the number of spatial dimensions of the model.
@@ IV Conclusions and Outlook
In summary, we have described the form of LIOMs in the Anderson model in Krylov space as generated by projectors on single sites. There are three types of behaviors – stretched exponential decay inside the localized phase (with geometric exponent $\gamma=1/2d$ set by real space dimension $d$), delocalization on real-space volume $V=\sqrt{D_{K}}$ states in the metallic/extended phase, and powerlaw decay at the critical point. Importantly, these results may be interpreted using effective disorder-free chains obtained from disorder averaged Krylov coefficients, which allows considerable further insight into the Anderson transition: e.g. the significance of Krylov dimerization as a mechanism of localizing operators despite having a gapless Liouvillian.
Our work may be extended to other aspects of the Anderson transition: for example, to perhaps enable the tracking of the mobility edge as a function of typical energy. It also sets up a framework for extending these results to interacting models where the coefficients grow with some power $b_{n}\sim n^{\delta}$ (with $\delta\leq 1$). If the dimerization decays like a powerlaw as in our model $\sim n^{-(1-\gamma)}$ then – since only the ratio of Lanczos coefficients matters for the zero-mode – the growth simply re-normalizes the power to be $\sim n^{-(1-\gamma-\delta)}$. Consequently, the faster the coefficients grow, i.e. the faster the operator growth, the slower the dimerization must decay in order for conserved quantities to exist. For spin systems debated to show MBL, such as the disordered Heisenberg model, the coefficients have been found to grow as $\sim n/\log{n}$ on average, with some dimerization [Trigueros2022_KrylovComplexityofMBL, Weisse2025_OperatorGrowthMBL]. The $\log$ correction would only be relevant at the critical power implying that the dimerization has to be constant for conserved quantities to exist (i.e. $\gamma=1$, its maximum allowed value). Additionally, it would be interesting to understand how the dimerization behaves in interacting integrable models such as the XXZ-chain, where there is typically square root growth ($\delta=0.5$) and thus only $1\geq\gamma\geq 0.5$ is allowed to host the conserved quantities we know to exist.